OpenStudy (anonymous):

Solve: −2(x − 5) = −4 A) 7 B) -7 C) -1/2 D) -3

hartnn (hartnn):

divide both sides by -2 so what is -4/-2 = ?

hartnn (hartnn):

what is 4/2 =?

OpenStudy (anonymous):

2

hartnn (hartnn):

then -4/-2 is also 2 now u have x-5=2 add 5 to both sides. what u get?

OpenStudy (anonymous):

would the answer be -7?

hartnn (hartnn):

x-5=2 x=5+2 =7 *positive* 7

OpenStudy (anonymous):

okay thanks :) @hartnn